OpenStudy (anonymous):

The surface area of a triangular pyramid is 408 m2. If all the faces of the pyramid are congruent, what is the area of one face of this pyramid? A. 51 m2 B. 68 m2 C. 136 m2 D. 102 m2

OpenStudy (anonymous):

How many faces do you think a triangular pyramid have?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Total Surface Area = # of Faces X Area of one Face

OpenStudy (anonymous):

You just have to divide that number to the surface area because the surface area is the total of all of each faces.

OpenStudy (anonymous):

408 = 4*X X=408/4 X = ?

OpenStudy (anonymous):

\[102m^{2}\]

OpenStudy (anonymous):

plus, they said all the faces are equal.

OpenStudy (anonymous):

Anyway I just gave you an idea how to solve this question.